OpenStudy (vshiroky):

There are three ways of solving systems of linear equations: graphing, substitution, and elimination. Post a problem, and show your step-by-step solution using one of the three methods. Now go and check out a classmates problem and solution. Are they correct? If not, are you able to point out their error so they can fix it? Dont forget to do a check. Put your solution into your original problem. Does it work

OpenStudy (godlovesme):

which of the methods is easy for u? @vshiroky

OpenStudy (vshiroky):

I think elimination is easiest for me. Sorry I'm doing 100 things at one time lol

OpenStudy (godlovesme):

it's fine but unfortunately i suck at elimination but it's fine so can u like make up an equation(with no fraction xD)

OpenStudy (vshiroky):

I am awful at making up equations I can prob do the work, but can make an equation lol

OpenStudy (godlovesme):

lolol so u want me to come up with one? :)

OpenStudy (vshiroky):

Please lol

OpenStudy (godlovesme):

xD k hang on

OpenStudy (godlovesme):

k here: 2x+5y=17 and 6x-5y=-9 lol i got from my textbook i'm soo creative :')

OpenStudy (vshiroky):

Haha awesome thank you!!

OpenStudy (godlovesme):

np :) so can u go ahead and try to solve and if u got stuck i'll help u :)

OpenStudy (vshiroky):

I will, but it might take me a while I'm a teacher and my prep is almost over lol

OpenStudy (godlovesme):

oh okay i feel so special helping a teacher xD jk :P

OpenStudy (vshiroky):

Lol yeah I'm a teacher and a student so this is my homework

OpenStudy (godlovesme):

cool i'm just a student ;( lol

OpenStudy (vshiroky):

I got (8,1/5) I think it's wrong

OpenStudy (vshiroky):

@Godlovesme

OpenStudy (godlovesme):

can u give me a sec please i started helping another person while u were gone sorry just 1 sec okay

OpenStudy (godlovesme):

@vshiroky i'm back

OpenStudy (vshiroky):

hey

OpenStudy (godlovesme):

hiii :)sooo how did u get that?

OpenStudy (vshiroky):

it's too much to type lol I still have kids in here

OpenStudy (godlovesme):

do u want me to wait a little

OpenStudy (vshiroky):

I didn't do it right I found online it's not the right answers

OpenStudy (godlovesme):

its fine we will work it out together :) do u see there is -5y and +5y ?

OpenStudy (godlovesme):

@vshiroky

OpenStudy (vshiroky):

I got it now... (1,3) Which book did you use just to make sure that we are not using the same book lol

OpenStudy (godlovesme):

me too but i used algebra 1

OpenStudy (vshiroky):

I made a problem can you see if you can solve it? 36x-6y=72 and 18x+6y=36

OpenStudy (godlovesme):

lemme try

OpenStudy (godlovesme):

i got (2,0)

OpenStudy (vshiroky):

Ok good me too :)

OpenStudy (godlovesme):

yay :D so what else do u need help with?

OpenStudy (vshiroky):

That's it for now lol

OpenStudy (vshiroky):

Thank you. I have to get going now

OpenStudy (godlovesme):

ok u r welcome :) take care lol